OpenStudy (anonymous):

m angle4=(5x+2) and m angle8=(x+50) what is the measure of angle 4 if the angles are corresponding angles?

Directrix (directrix):

These angles could be corresponding angles of parallel lines cut by a transversal or perhaps corresponding angles of congruent polygons. That is irrelevant given that the angles are equal in measure. In fact, giving the angles as equal in measure precludes the necessity of writing that the angles are corresponding. 5x + 2 = x + 50 4x = 48 x = 12 m angle4=(5x+2) = 5(12) +2 = 62.